OpenStudy (anonymous):

A turntable rotates with a constant angular acceleration 2.30rad/s^2 . After a time interval 3.75s the wheel has rotated through an angle of 59.0rad . What was the angular velocity of the wheel at the beginning of the time interval 3.75s ?

OpenStudy (anonymous):

L 2 u p v o t e

OpenStudy (anonymous):

Do..u have answer key with u ....@michelle092294

OpenStudy (anonymous):

no i dont:( this is online assignment

OpenStudy (anonymous):

oh i got it:)

OpenStudy (anonymous):

wait its still wrong -_-

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\alpha = \frac{ d \omega }{ dt }\]

OpenStudy (anonymous):

angular accl = change in angular velocity

OpenStudy (anonymous):

\[2.3 * 3.75 = \delta \omega\]

OpenStudy (anonymous):

\[\omega f - \omega i = 8. 625\]

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

lol....not finished Yet....

OpenStudy (anonymous):

\[\omega i = \omega f - 8.625\]

OpenStudy (anonymous):

\[\omega f = 56 / 3.75 = 15.73\]

OpenStudy (anonymous):

wrong.

OpenStudy (anonymous):

yeah. its wrong . i try that answer awhile ago -_-

OpenStudy (anonymous):

\[\omega i = 15.73 - 8.625 = 7.10\]

OpenStudy (anonymous):

lol.....Guys wait..) let me Finish

OpenStudy (anonymous):

more wrong.

OpenStudy (anonymous):

sure please finish

OpenStudy (anonymous):

Do u have options ..)

OpenStudy (anonymous):

nope. i dont have

OpenStudy (anonymous):

Hm....... My appro Answer is 7.10

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

wrong -_-

OpenStudy (anonymous):

hm.....)

OpenStudy (anonymous):

help me :(

OpenStudy (paxpolaris):

\[\large 59=\omega_i \cdot 3.75+\frac12 \cdot 2.30\cdot (3.75)^2\] \[\Large \omega_i= {59 \over 3.75} - \frac 12 \cdot 2.30 \cdot3.75\]

OpenStudy (anonymous):

thanks :)

OpenStudy (paxpolaris):

\[\approx 11.4 radians /second\]

OpenStudy (anonymous):

got it :)